5/3x+1/3x=5x+14/3+5/3x
WHat is the solution for x?
THank you

Answer by jojo14344(1513)

5/3x+1/3x=5x+14/3+5/3x
becomes,
5x-1/3x +5/3x-5/3x= -14/3
5x-1/3x= -14/3
14/3x = -14/3
14x = -14
x = -1
This will be the answer supposedly "x" is a numerator, not a denominator on the equation. Thank you.